# Probing Mechanical Reasoning in Large Vision Language Models

### v3

@@ 5 Conclusion
In this paper, we presented the first comprehensive assessment of VLMs’ mechanical reasoning abilities. Drawing on experimental paradigms from cognitive science literature, our findings reveal that current VLMs fail to fully acquire the necessary competencies for mechanical reasoning across various domains. Furthermore, the observation that performance on certain tasks does not improve with increased model size underscores potential fundamental limitations in existing model architectures. In line with previous research, we specifically highlighted model-based reasoning—the ability to mentally simulate dynamic relations within physical systems. However, given that mechanical reasoning is a high-level cognitive ability reliant on a range of basic cognitive processes, the precise mechanisms underlying these limitations remain unclear. Further exploration of these mechanisms could provide crucial insights into the constraints of current AI models and inform the development of more capable architectures for mechanical reasoning tasks.
@@
Appendix A Detailed Analysis of Model Performance on Each Domains of Mechanical Reasoning
@@ A.1 System Stability
VLMs performance in intuitively evaluating system stability is not ideal, however here we observe some interesting phenomena. Models excel at identifying objects in images, and they not only understand what the objects are, the models also successfully recognize their mechanical states as well. Also, the models could effectively connect mechanical descriptions with the corresponding mechanical scenarios (Figure 7). In Experiments 7B and 7F, the models could recognize the two chairs and the two bottles, and their mechanical situations; however, the models still failed to provide the correct answers in the experiments. For instance, in Experiment 7B, the model explains, ”the stool on the right is more likely to tip over when an active child sits on it because its legs are splayed at an angle, making the base wider and potentially more stable under normal circumstances but also adding a tipping hazard due to the non-vertical configuration of the legs”. The model correctly notices that it’s the angle of the leg that matters for the stability of the system. However, it reasons completely the opposite way to correct answers. When the leg’s angle wider, it’s actually more stable. It’s a very intuitive physical problem for humans but the models fail, even though they still demonstrate step-by-step reasoning abilities in this case. In Experiment 7F, the model is correct that the bottle on the bottom ”has a larger area of contact with the surface, creating more friction, which helps prevent rolling”. However, the model fails to realize one bottle is standing, and one bottle is rolling, and sliding friction and rolling friction are completely different. In contrast, humans can easily solve these problems intuitively.
@@ A.2 Pulley Systems
VLMs perform poorly in recognizing movable pulley systems. In Experiment 8A, the image includes a standard single movable pulley system and a standard single fixed pulley system. The question ”Which system requires less effort?” is essentially asking whether the model can correctly select the movable pulley. Clearly, the model failed in its selection, as it straightforwardly provided an incorrect answer in its explanation. VLMs also struggle in predicting whether a suspended weight is being lifted or lowered through a pulley system. Experiments 8B-F either directly or indirectly reflect this issue, with Experiment 8D being the most direct and concise. In Experiment D, the weight is directly attached to the movable pulley, and by pulling the other end of the rope, the pulley and the weight are lifted. However, the model’s response was the exact opposite of the correct answer. In its explanation, the model seemed to imply that the pulley was not fixed (though it did not explicitly state that it was a movable pulley), and the physics it provided was entirely incorrect. Therefore, we can hypothesize that the model’s poor performance in predicting the weight’s movement may be due to its limited ability to recognize movable pulleys. However, the specific reasons require further experiments to be analyzed in detail.
The above two issues confirm that VLMs still have limitations in recognizing pulley systems, whereas for individuals with some mechanical experience, identifying simple pulley systems through basic diagrams is not difficult (especially in the case of Experiments 8A and 8D).
@@ A.4 Seesaw-like systems and Leverage Principle
We observe diverse behaviors in VLMs on solving seesaw-like mechanical systems and applying leverage principle (Figure 10). Experiments 10A, 10C, 10D, and 10F involve relatively simple scene-based question answering and basic extrapolation. Experiments 10A and 10D are straightforward question-answer tasks. As long as the model identifies the key information in the images—such as the length of the wrench in Experiment 10A or the unequal torques in Experiment 10D—it can easily solve the problems. Experiments 10C and 10F involve simple extrapolations based on scene information, like understanding how changes in the length of the resistance arm affect force in Experiment 10C or predicting the outcome when the weights on a balance scale are unequal in Experiment 10F. It is evident that the model performed exceptionally well in these experiments, providing both correct answers and detailed explanations. However, in Experiments 10B and 10E, VLMs failed to give the correct answers. We suspect that VLMs struggle when the mechanical reasoning processes, particularly applying the leverage principle and solving seesaw-like systems, require multiple steps. For example, in Experiment 10B, while VLMs recognize it is a seesaw-like system, they fail to understand how to operate within this system: sliding backward would give more leverage. This reasoning process may require backtracking through several steps: first, identifying the seesaw-like system; second, understanding how to apply leverage; third, realizing that sliding backward increases leverage; fourth, recognizing that more leverage is needed to level the system; and finally, concluding that sliding backward is necessary. The same issue arises in Experiment 5E.
@@ A.5 Inertia and Motion
Experiment 11E involves scene-based question. The model not only identified that the cart was not stationary, but it also further analyzed that if the object suspended on the cart were stationary, it would fall vertically due to gravity. However, the depicted scene likely indicates that the cart is accelerating forward. Experiments 11A, 11B, and 11D, on the other hand, are cases of simple extrapolation. They are categorized as ”simple extrapolation” because the scenarios are straightforward and involve a single, clear change. For example, in Experiment 11A, the change involves the cart suddenly stopping—a relatively simple scene (involving only two objects, the drink and the cart, with straight-line motion and vertical force equilibrium). Similar patterns apply to Experiments 11B and 11D.
Experiments 11C and 11F, however, involve more complex reasoning related to inertia and motion prediction. Experiment 11F is a derivative question from Experiment 11E, asking what would happen if the cart suddenly stopped. The suspended object would continue moving forward due to inertia. Although the scene in Experiment 11F is similar to Experiment 11E, and the question is similar to that in Experiment 11A, it involves more objects (the pulling rope, the cart, and the pulled object), more physical principles, and a sudden change in force (as the rope loses tension). The model clearly struggled with this problem, providing an incorrect answer and an explanation that did not meet expectations. Notably, Experiment 11C asks a common-sense question based on pendulum motion. The image in Experiment 11C marks three points in the half-arc of the pendulum’s swing: the highest point, the quarter-arc point, and the lowest point. This information is also specified in the prompt. However, the model gave an incorrect answer, and its explanation was confused. The model completely misunderstood the designated points in the diagram and in the prompt, mistaking the first and third points as the left and right highest points of the pendulum’s motion and the second point as the lowest point. One possible reason for this error is that the model may have mistaken the half-arc pendulum motion in Experiment 11C for a full pendulum motion. A deeper explanation might suggest that the model has issues with perceptual constancy. The image in Experiment 11C might have been misinterpreted as a 3D scene, and this visual misperception could be the root of the error. This hypothesis requires further experimentation to be confirmed.
Figure 12: VLMs performance on reasoning about fluid systems in MechBench
A.6 Fluid Mechanics
@@
Experiments 12B and 12C highlighted the model’s weaker inference abilities. In Experiment 12B, the hole in the bucket was positioned at the middle of the bucket’s wall. Clearly, once the water level drops to the level of the hole, water will stop flowing out, meaning some water will remain in the bucket. However, the model failed to capture this crucial detail about the hole’s location, leading to an incorrect answer. Similarly, in Experiment 12C, although the model provided the correct answer, its explanation was incorrect.
It is worth noting that the model excelled in image comprehension and detail recognition, especially in Experiments 12D and 12F. In Experiment 12D, the measuring cup had a narrow base and a wider top, meaning the scale markings could not be evenly spaced. The model successfully captured these details, including the design of the measuring cup, the distribution of the scale, and the corresponding numerical values and units, offering a detailed explanation. The same can be said for Experiment 12F. The model accurately identified that the balloon was fully inflated with gas and incorporated the concepts of gas and liquid density to explain the current state and predict the next step of the experiment.

### v4

@@ 5 Conclusion
In this paper, we presented the first comprehensive assessment of VLMs’ mechanical reasoning abilities. Drawing on experimental paradigms from cognitive science literature, our findings reveal that current VLMs fail to fully acquire the necessary competencies for mechanical reasoning across various domains. Furthermore, the observation that performance on certain tasks does not improve with increased model size underscores potential fundamental limitations in existing model architectures. In line with previous research, we specifically highlighted model-based reasoning—the ability to mentally simulate dynamic relations within physical systems. However, given that mechanical reasoning is a high-level cognitive ability reliant on a range of basic cognitive processes (Kaiser et al., 1986; Hegarty et al., 1988; Allen, 2021), the precise mechanisms underlying these limitations remain unclear. Further exploration of these mechanisms could provide crucial insights into the constraints of current AI models and inform the development of more capable architectures for mechanical reasoning tasks.
@@
Appendix A Detailed Analysis of Model Performance on Each Domain of Mechanical Reasoning
@@ A.1 System Stability
Models excel at identifying objects in images, and they not only understand what the objects are, the models also successfully recognize their mechanical states as well. Also, the models could effectively connect mechanical descriptions with the corresponding mechanical scenarios (Figure 6). In Experiments 6B and 6F, the models could recognize the two chairs and the two bottles, and their mechanical situations; however, the models still failed to provide the correct answers in the experiments. For instance, in Experiment 6B, the model explains, ”the stool on the right is more likely to tip over when an active child sits on it because its legs are splayed at an angle, making the base wider and potentially more stable under normal circumstances but also adding a tipping hazard due to the non-vertical configuration of the legs”. The model correctly notices that it’s the angle of the leg that matters for the stability of the system. However, it reasons completely the opposite way to correct answers. When the leg’s angle wider, it’s actually more stable. It’s an intuitive physical rule problem for humans but the models fail, even though they still demonstrate step-by-step reasoning abilities in this case. In Experiment 6F, the model is correct that the bottle on the bottom ”has a larger area of contact with the surface, creating more friction, which helps prevent rolling”. However, the model fails to realize one bottle is standing, and one bottle is rolling, and sliding friction and rolling friction are completely different.
@@ A.2 Pulley Systems
VLMs perform poorly in recognizing movable pulley systems. In Experiment 7A, the image includes a standard single movable pulley system and a standard single fixed pulley system. The question ”Which system requires less effort?” is essentially asking whether the model can correctly select the movable pulley. Clearly, the model failed in its selection, as it straightforwardly provided an incorrect answer in its explanation. VLMs also struggle in predicting whether a suspended weight is being lifted or lowered through a pulley system. Experiments 7B-F either directly or indirectly reflect this issue, with Experiment 7D being the most direct and concise. In Experiment D, the weight is directly attached to the movable pulley, and by pulling the other end of the rope, the pulley and the weight are lifted. However, the model’s response was the exact opposite of the correct answer. In its explanation, the model seemed to imply that the pulley was not fixed (though it did not explicitly state that it was a movable pulley), and the physics it provided was entirely incorrect. Therefore, we can hypothesize that the model’s poor performance in predicting the weight’s movement may be due to its limited ability to recognize movable pulleys. However, the specific reasons require further experiments to be analyzed in detail.
The above two issues confirm that VLMs still have limitations in recognizing pulley systems, whereas for individuals with some mechanical experience, identifying simple pulley systems through basic diagrams is not difficult (especially in the case of Experiments 7A and 7D).
@@ A.4 Seesaw-like systems and Leverage Principle
We observe diverse behaviors in VLMs on solving seesaw-like mechanical systems and applying leverage principle (Figure 9). Experiments 9A, 9C, 9D, and 9F involve relatively simple scene-based question answering and basic extrapolation. Experiments 9A and 9D are straightforward question-answer tasks. As long as the model identifies the key information in the images—such as the length of the wrench in Experiment 9A or the unequal torques in Experiment 9D—it can easily solve the problems. Experiments 9C and 9F involve simple extrapolations based on scene information, like understanding how changes in the length of the resistance arm affect force in Experiment 9C or predicting the outcome when the weights on a balance scale are unequal in Experiment 9F. It is evident that the model performed exceptionally well in these experiments, providing both correct answers and detailed explanations. However, in Experiments 9B and 9E, VLMs failed to give the correct answers. We suspect that VLMs struggle when the mechanical reasoning processes, particularly applying the leverage principle and solving seesaw-like systems, require multiple steps. For example, in Experiment 9B, while VLMs recognize it is a seesaw-like system, they fail to understand how to operate within this system: sliding backward would give more leverage. This reasoning process may require backtracking through several steps: first, identifying the seesaw-like system; second, understanding how to apply leverage; third, realizing that sliding backward increases leverage; fourth, recognizing that more leverage is needed to level the system; and finally, concluding that sliding backward is necessary. The same issue arises in Experiment 5E.
@@ A.5 Inertia and Motion
Experiment 10E involves scene-based question. The model not only identified that the cart was not stationary, but it also further analyzed that if the object suspended on the cart were stationary, it would fall vertically due to gravity. However, the depicted scene likely indicates that the cart is accelerating forward. Experiments 10A, 10B, and 10D, on the other hand, are cases of simple extrapolation. They are categorized as ”simple extrapolation” because the scenarios are straightforward and involve a single, clear change. For example, in Experiment 10A, the change involves the cart suddenly stopping—a relatively simple scene (involving only two objects, the drink and the cart, with straight-line motion and vertical force equilibrium). Similar patterns apply to Experiments 10B and 10D.
Experiments 10C and 10F, however, involve more complex reasoning related to inertia and motion prediction. Experiment 10F is a derivative question from Experiment 10E, asking what would happen if the cart suddenly stopped. The suspended object would continue moving forward due to inertia. Although the scene in Experiment 10F is similar to Experiment 10E, and the question is similar to that in Experiment 10A, it involves more objects (the pulling rope, the cart, and the pulled object), more physical principles, and a sudden change in force (as the rope loses tension). The model clearly struggled with this problem, providing an incorrect answer and an explanation that did not meet expectations. Notably, Experiment 10C asks a common-sense question based on pendulum motion. The image in Experiment 10C marks three points in the half-arc of the pendulum’s swing: the highest point, the quarter-arc point, and the lowest point. This information is also specified in the prompt. However, the model gave an incorrect answer, and its explanation was confused. The model completely misunderstood the designated points in the diagram and in the prompt, mistaking the first and third points as the left and right highest points of the pendulum’s motion and the second point as the lowest point. One possible reason for this error is that the model may have mistaken the half-arc pendulum motion in Experiment 10C for a full pendulum motion. A deeper explanation might suggest that the model has issues with perceptual constancy. The image in Experiment 10C might have been misinterpreted as a 3D scene, and this visual misperception could be the root of the error. This hypothesis requires further experimentation to be confirmed.
Figure 11: VLMs performance on reasoning about Liquid Mechanics in MechBench
A.6 Liquid Mechanics
@@
Experiments 11B and 11C highlighted the model’s weaker inference abilities. In Experiment 11B, the hole in the bucket was positioned at the middle of the bucket’s wall. Clearly, once the water level drops to the level of the hole, water will stop flowing out, meaning some water will remain in the bucket. However, the model failed to capture this crucial detail about the hole’s location, leading to an incorrect answer. Similarly, in Experiment 11C, although the model provided the correct answer, its explanation was incorrect.
It is worth noting that the model excelled in image comprehension and detail recognition, especially in Experiments 11D and 11F. In Experiment 11D, the measuring cup had a narrow base and a wider top, meaning the scale markings could not be evenly spaced. The model successfully captured these details, including the design of the measuring cup, the distribution of the scale, and the corresponding numerical values and units, offering a detailed explanation. The same can be said for Experiment 11F. The model accurately identified that the balloon was fully inflated with gas and incorporated the concepts of gas and liquid density to explain the current state and predict the next step of the experiment.